After a late-winter freeze in Florida, the supply curve for early-season blueberries was estimated as: QS = - 500 + 5 000 P The demand curve for blueberries is:

QD = 19 000 – 1 500 P P : measured in dollars, Q : measured in pints

3) How much are consumers willing to pay for blueberries when this many pints of blueberries are sold? What is the maximum value of time consumers would be willing to spend waiting in line for blueberries?


4) Suppose that sellers of blueberries must spend $0.50 per pint to pick blueberries. The sellers try to evade the price ceiling by selling only U-pick blueberries. Consumers value U-pick pints of blueberries at 0.50 per pint less than picked blueberries because of the opportunity cost of picking time. If all blueberries sold are U-pick blueberries, how many pints of blueberries are sold?


4) What is the equilibrium price of U-pick blueberries?

5) Is the outcome in 4) more efficient than the outcome in 2)?


1
Expert's answer
2021-12-05T18:54:18-0500

Solution:

3.). At equilibrium: QD = QS

19,000 – 1,500P = -500 + 5,000P

19,000 + 500 = 5000P + 1500P

19,500 = 6500P

P = 0.3

The maximum value of time consumers would be willing to spend waiting in line for blueberries = 0.3                                                                                                                                                                                                                                                          

 

Derive the quantity:

QD = 19,000 – 1,500P = 19,000 – 1,500(0.3) = 19,000 – 450 = 18,550

 

4.).  The equilibrium price of brick of U-pick blueberries = 0.3                                                                                                                                                                                                                                                                                                                                                                            The option is more efficient than the outcome in   outcome 2